OpenStudy (anonymous):

Which is a zero of f(x) = x2 - 15x + 54?

OpenStudy (anonymous):

thats so easy

OpenStudy (***[isuru]***):

u r wrong.....

OpenStudy (anonymous):

9 5 3 15 are one of the anwsers.?

OpenStudy (***[isuru]***):

u have miss the factorization x2 - 15x + 54 = x2 - 6x - 9x +54 = x( x -6) -9 ( x -6) = (x-6)(x-9) this is the correct factorization

OpenStudy (anonymous):

x2-15X + 54 = (X - 6)(X -9)

OpenStudy (***[isuru]***):

so.. x = 6 or x = 9

OpenStudy (anonymous):

I said that

OpenStudy (***[isuru]***):

i think now u have an answer... juliogomez

OpenStudy (anonymous):

Everyone has a different ways of doing their work.

OpenStudy (anonymous):

The sun is about 1.5 × 108 kilometers from Earth. If light travels about 3 × 105 kilometers per second, about how many seconds does it take light from the sun to reach Earth? help me :(

OpenStudy (***[isuru]***):

i was joking.. if u were on the correct path.. then i appreciate ur work ... i interrupted 'cause i thought u may direct him to the wrong path

OpenStudy (***[isuru]***):

keep up ur good work :)

OpenStudy (anonymous):

500 seconds

OpenStudy (anonymous):

thanks (;

OpenStudy (anonymous):

Your welcome :)